A lightning bolt transfers 30 C of charge to Earth through an average potential difference of 40 MV.

Unit 3: Prelecture Problems Problems: Electric Potential & Capacitor Deadline: 100% until Sunday, January 27 at 11:59 PM Cumulative Problem 4 A lightning bolt transfers 30 C of charge to Earth through an average potential difference of 40 MV. 1) How much energy is dissipated in the bolt? (Express your answer to two significant figures.) GJ Submit You currently have 0 submissions for this question. Only 10 submission are allowed. You can make 10 more submissions for this question.
Answer #1

Energy dissipated in the bolt will be given by:

U = q*dV

q = charge transferred by lightning bolt = 30 C

dV = potential difference = 40 MV = 40*10^6 V

So Using these values:

U = q*dV = 30 C*(40*10^6 V)

U = 1.2*10^9 J

U = 1.2 GJ

Please Upvote.

Comment below if you have any query.